OpenStudy (anonymous):

The density (mass divided by volume) of water is 1.00g/cm3 . What is this value in kilograms per cubic meter?

OpenStudy (akashdeepdeb):

To solve these types of questions... Convert the units to the required units... You are given 1 gram How many kilograms is 1 gram? [ 1/1000 kilograms right?] Convert 1cm^3 to m^3 [How much will it be?] 1cm^3 = 1/1000000 m^3 So \[1 g/cm^3 = \frac{ 1/1000 kg }{ 1/1000000m^3 } = 1000kg/m^3\] Understood? :) To people who have helped you, "Click the Best Response Button" :)

OpenStudy (anonymous):

Thanks, I understand the math, I just am not good at figuring out the conversions. Not sure where to start or to look ever.

OpenStudy (akashdeepdeb):

But you got this right? :D

OpenStudy (anonymous):

Yes, thanks! :)

OpenStudy (akashdeepdeb):

:)